b.  RM 2500 is invested for 4 years 9 months with interest rates of 8% compounded monthly for the first two years and 6.5% compounded annually for the rest of the period. Find the accumulated amount and interest earned.



1
Expert's answer
2021-06-29T03:46:06-0400

Solution.

1) first two years = first 24 months

From the formula


"A_n=A_0(1+\frac{p}{100})^n"

we will have


"A_n=2500(1+\frac{\frac{8}{12}}{100})^{24}=2932."

2) next period=33 months

From the formula


"A_n=A_0(1+\frac{p}{100})^n"

we will have


"A_n=2932(1+\frac{6.5}{100})^{\frac{33}{12}}=3486."

3) The accumulated amount is equal 3486 RM.

Interest earned is 3486-2500= 986 RM.